# A kernel conditional two-sample test

### v1

@@ 6 Conclusion
We rigorously introduce a framework for testing conditional distributions with local guarantees. The framework is broadly applicable and provides a general recipe for turning learning guarantees into statistical tests. We instantiate it for conditional two-sample testing with KRR, enabling the comparison of not only expectations but also arbitrary moments and full conditional distributions. The guarantees are based on a generalization we show of popular confidence bounds for KRR and regularized least squares to accommodate UBD kernels. To our knowledge, this is the first generalization to support such a broad class of possibly non-trace-class kernels. We present a complete pipeline from conceptual foundations and theoretical analysis to practical implementation, and validate the approach through numerical experiments. Our framework offers a principled foundation for developing powerful, flexible, and theoretically grounded conditional tests, laying a basis for a wide range of future applications or theoretical investigations.
@@ Acknowledgment
Friedrich Solowjow is supported by the German federal state of North Rhine-Westphalia through the KI-Starter grant.
Appendix A Tutorials
The goal of this section is to demonstrate how to apply in practice the tools described in the main text. We begin by explaining how we can modify the two-sample hypothesis (5) to compare not only expectations, but also other properties of conditional distributions, such as higher moments or the full conditional distributions themselves. We then provide algorithms summarizing the test and the bootstrap approach (cf. Appendix B). Finally, we showcase how to use the method to detect changes in the dynamics of a linear dynamical system.
A.1 Comparing other properties of conditional distributions
Our test can be used to compare properties of conditional distributions other than only their means. For the sake of discussion, assume we are interested in testing not $H_{0}$ as in (5), but $H_{0,m}$ defined as
$H_{0,m}(x,p_{1},p_{2}):\text{the first $m$ moments of $p_{1}(\cdot,x)$ and $p_{2}(\cdot,x)$ coincide},$ (12)
where $m\in{\mathbb{N}_{>0}}\cup\{\infty\}$, with the convention that $H_{0,\infty}(x,p_{1},p_{2})$ is simply “$p_{1}(\cdot,x)=p_{2}(\cdot,x)$”. We describe below how to achieve testing of $H_{0,m}$.
Transforming the data
The key idea we leverage is that Theorem 4.4 holds for all [^5] In all of this section, we omit technicalities to keep the focus on applications. This means that we do not recall technical assumptions such as separability or measurability and refer to the main text or the other appendices for the technically-precise statements. Hilbert output spaces $\mathcal{G}$. In particular, this allows applying the test not directly on the measurements, but rather on a function of the measurements (say, $\Phi_{m}$) that we pick appropriately such that $H_{0,m}$ is equivalent to $H_{0}$ on the transformed output. Let us introduce this concept on the following example.
Example A.1 (Comparing means and variances).
Let $\mathcal{Z}$ be a subset of $\mathbb{R}$, which we call the measurement set, and assume we have two Markov kernels $p_{1}$ and $p_{2}$ from $\mathcal{X}$ to $\mathcal{Z}$, respectively. Also assume that $p_{1}$ and $p_{2}$ have well-defined variances (e.g., it suffices that $\mathcal{Z}$ is bounded) and introduce for $i\in\{1,2\}$ the conditional first two moments:
$\mu^{(i)}_{1}(x)=\mathbb{E}_{Z\sim p_{i}(\cdot,x)}[Z],\quad\text{and}\quad\mu^{(i)}_{2}(x)=\mathbb{E}_{Z\sim p_{i}(\cdot,x)}[Z^{2}]$
for all $x\in\mathcal{X}$. It is clear that $H_{0,2}(x,p_{1},p_{2})$ is equivalent to the statement that $\begin{pmatrix}\mu^{(1)}_{1}(x)&\mu^{(1)}_{2}(x)\end{pmatrix}=\begin{pmatrix}\mu^{(2)}_{1}(x)&\mu^{(2)}_{2}(x)\end{pmatrix}$. Denote now as $q_{i}(\cdot,x)$ the distribution of the variable $Y=\begin{pmatrix}Z&Z^{2}\end{pmatrix}=:\Phi_{2}(Z)$, where $Z\sim p_{i}(\cdot,x)$, for each $i\in\{1,2\}$. This corresponds to defining $q_{i}$ as $q_{i}(\cdot,x)=p_{i}(\Phi_{2}^{-1}(\cdot),x)$ (this construction is also known as the pushforward of $p_{i}$ through $\Phi_{2}$). We have
$\mathbb{E}(q_{i})(x)$ $=\int_{\mathbb{R}^{2}}yq_{i}(\mathrm{d}y,x)$
$=\begin{pmatrix}\int_{\mathbb{R}}zp_{i}(\mathrm{d}z,x)\\ \int_{\mathbb{R}}z^{2}p_{i}(\mathrm{d}z,x)\end{pmatrix}$
$=\begin{pmatrix}\mu^{(i)}_{1}(x)\\ \mu^{(i)}_{2}(x)\end{pmatrix},$
where the second equality results from standard calculation rules for pushforward measures. In other words, $H_{0,2}(x,p_{1},p_{2})$ holds if, and only if, $H_{0}(x,q_{1},q_{2})$ does.
A natural next step is to try and test for $H_{0}(x,q_{1},q_{2})$ from data. To this end, let $D^{(1)}=(X^{(1)}_{n},Z^{(1)}_{n})_{n\in{\mathbb{N}_{>0}}}$ and $D^{(2)}=(X^{(2)}_{n},Z^{(2)}_{n})_{n\in{\mathbb{N}_{>0}}}$ be processes of transition pairs of $p_{1}$ and $p_{2}$, respectively. Crucially, for all $i\in\{1,2\}$, we are able to create a process of transition pairs of $q_{i}$ from $D^{(i)}$ by simply taking $D^{(i)}_{\Phi_{2}}=(X^{(i)}_{n},\Phi_{2}(Z^{(i)}_{n}))_{n\in{\mathbb{N}_{>0}}}$. Theorem 4.2 thus guarantees that that any learning algorithm on $D^{(i)}_{\Phi_{2}}$ with confidence bounds yields a test of $H_{0}(x,p_{1},p_{2})$. This is for instance the case for KRR. One should bear in mind, however, that introducing the data transformation $z\mapsto\Phi_{2}(z)$ made the output $2$-dimensional, requiring vector-valued KRR, for which an accessible introduction to the theory and implementation can be found in Alvarez et al. [37].
Choice of kernel on the input set
Example A.1 allows any kernel function $K:\mathcal{X}\times\mathcal{X}\to\mathbb{R}^{2\times 2}$ for the resulting KRR (as long as the assumptions of Theorem 4.4 hold, to have guarantees). In practice, however, one often makes the simplifying choice $K(x,x^{\prime})=k(x,x^{\prime})\cdot\mathrm{id}_{\mathbb{R}^{2}}\approx k(x,x^{\prime})I_{2}$, where $I_{2}$ is the identity matrix in $\mathbb{R}^{2\times 2}$ and $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is a scalar kernel. While there is no a priori reason for this as it is ultimately a modeling choice, an argument is that the kernel $k$ should capture how the conditional distributions $p_{1}$ and $p_{2}$ vary as functions of the covariate, and the introduction of a vector-valued output is only a trick we use to reformulate our hypothesis in the form (5). With this choice for $K$, the expression of the estimator (4) of KRR simplifies to (up to transposition to obtain a column vector)
$f_{D,\lambda}(x)=k(x,X)(k(X,X)+\lambda I_{n})^{-1}Y,$ (13)
where $D=((X_{1},Z_{1}),\dots,(X_{n},Z_{n}))$, $k(x,X)$ and $k(X,X)$ are respectively the usual kernel row vector and matrix (see also (8)), and $Y=(\Phi_{2}(Z_{1}),\dots,\Phi_{2}(Z_{n}))^{\top}\in\mathbb{R}^{n\times 2}$.
Generalization via the kernel trick
The preceding construction is in no way specific to the mean and variance and can easily be generalized to compare other aspects of the distributions $p_{1}(\cdot,x)$ and $p_{2}(\cdot,x)$. An inspection of the method reveals that the main degree of freedom lies in the choice of the map $\Phi$ (which we took as $\Phi_{2}$ in Example A.1). We simply need to pick $\Phi$ such that our statement of interest holds if, and only if, $\mathbb{E}[q_{1}](x)=\mathbb{E}[q_{2}](x)$, with $q_{i}(\cdot,x)=p_{i}(\Phi^{-1}(\cdot),x)$, and to apply the transformation $z\mapsto\Phi(z)$ to the output data. In other words, we are doing feature engineering so that the expectation of our transformed output vector captures exactly the property we are interested in, and the map $\Phi$ is simply the “feature extractor”. Conversely, to any map $\Phi$ corresponds the hypothesis
$H_{0,\Phi}(x,p_{1},p_{2}):\quad\mathbb{E}[q_{1}^{\Phi}](x)=\mathbb{E}[q_{2}^{\Phi}](x),$
where $q_{i}^{\Phi}$ is the pushforward of $p_{i}$ through $\Phi$. We thus see that it is possible to capture complex statements about Markov kernels from $\mathcal{X}$ to $\mathcal{Z}$ in $H_{0}$ as defined in (5), by defining $\mathcal{G}$ and an embedding therein appropriately.
This raises the immediate question of computational aspects. Indeed, if $\Phi:\mathcal{X}\to\mathcal{G}$ takes values in a high dimensional space (say, we wish to compare the first $m$ moments, where $m\in{\mathbb{N}_{>0}}$ is large), then evaluating the difference $\lVert f_{D_{1},\lambda_{1}}(x)-f_{D_{2},\lambda_{2}}(x)\rVert_{\mathcal{G}}$ for the test (7) can turn out to be expensive, as it requires evaluating the scalar products $\langle\Phi(Z^{(i)}_{r}),\Phi(Z^{(j)}_{s})\rangle_{\mathcal{G}}$, where $i,j\in\{1,2\}$, $r\in\left[n_{i}\right]$, and $s\in\left[n_{j}\right]$, with $n_{1}=\mathrm{len}(D_{1})$ and $n_{2}=\mathrm{len}(D_{2})$. Further, in the case of high-order moments, computations with high powers can introduce numerical instability. We can address this by performing the kernel trick on the measurement set $\mathcal{Z}$. Specifically, introduce another kernel function $\kappa:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$, let $\mathcal{H}_{\kappa}$ be its RKHS, and take $K=k\cdot\mathrm{id}_{\mathcal{H}_{\kappa}}$, where $k$ is again a scalar kernel on $\mathcal{X}$. Then, the estimator (4) has the simple form
@@ A.1 Comparing other properties of conditional distributions
$f_{D,\lambda}(x)=k(x,X)(k(X,X)+\lambda I_{n})^{-1}\begin{pmatrix}\kappa(\cdot,Z_{1})\\ \vdots\\ \kappa(\cdot,Z_{n})\end{pmatrix},$ (14)
and the scalar products $\langle\kappa(\cdot,Z_{r}),\kappa(\cdot,Z_{s})\rangle_{\mathcal{H}_{\kappa}}$ appearing when computing the test statistic are simply $\kappa(Z_{r},Z_{s})$, by the reproducing property for the kernel $\kappa$.
The main remaining question is: what property of $p_{1}$ and $p_{2}$ is being compared when proceeding as such with a kernel $\kappa$? While answering it for an arbitrary kernel is nontrivial, the answer is known for many common kernels [26]. For instance, taking for $\kappa$ the inhomogeneous polynomial kernel of degree $m\in{\mathbb{N}_{>0}}$, that is, $\kappa(z,z^{\prime})=(1+\langle z,z^{\prime}\rangle_{\mathcal{Z}})^{m}$, enables comparing the first $m$ moments (here, $\mathcal{Z}\subset\mathbb{R}^{d})$. In contrast, taking any characteristic kernel [38] (such as the Gaussian or Matérn kernels) enables comparing the full conditional distributions.
Finally, we point the reader interested into going further on this topic to Muandet et al. [26, Chapter 4] and Park and Muandet [39], as the concurrent choices $\Phi(z)=\kappa(\cdot,z)$ and $K=k\cdot\mathrm{id}_{\mathcal{H}_{\kappa}}$ make KRR recover exactly the framework of kernel conditional mean embeddings: the space $\mathcal{H}_{K}$ is (canonically isomorphic [40] to) the space where the conditional mean embedding operator of [26, Equation (4.4)] lies, and the estimator (14) is precisely the usual empirical estimator of the conditional mean embedding.
Numerical study

### v2

@@ 6 Conclusion
We rigorously introduce a framework for testing conditional distributions with local guarantees. The framework is broadly applicable and provides a general recipe for turning learning-theoretic bounds into statistical tests. We instantiate it for conditional two-sample testing with KRR, enabling the comparison of not only expectations but other functionals such as arbitrary moments or full conditional distributions. The guarantees are based on a generalization we show of popular confidence bounds for KRR and regularized least squares to accommodate UBD kernels. To our knowledge, this is the first generalization to support such a broad class of possibly non-trace-class kernels. We present a complete pipeline from conceptual foundations and theoretical analysis to an algorithmic implementation, and validate the approach through numerical experiments. Our framework offers a principled foundation for developing powerful, flexible, and theoretically grounded conditional tests, laying a basis for a wide range of future applications or theoretical investigations.
@@ Acknowledgment
Friedrich Solowjow is supported by the German federal state of North Rhine-Westphalia through the KI-Starter grant. Christian Fiedler acknowledges support from DFG Project FO 767/10-2 (eBer-24-32734) "Implicit Bias in Adversarial Training".
Appendix A Elaboration on functional tests
The goal of this section is to expand on the observation of Section 4.3 that Theorem 4.2 provides a test for the generalized hypothesis $H_{0}$ defined in (12) for any functional $\mathcal{F}$ satisfying Assumption 4.5. Specifically, we show and explain the following:
  1. 1.
    There are many functionals $\mathcal{F}$ that satisfy Assumption 4.5.
  2. 2.
    There is a general catalog of representation maps $\Phi_{\mathcal{F}}$ of functionals capturing specific properties of interest of the Markov kernels — such as moments or full distributions. The catalog consists of kernel functions, and the properties are the ones captured by the corresponding kernel mean embeddings. This is summarized in Observation A.2.
  3. 3.
    Using KRR with a diagonal kernel $K=k\cdot\mathrm{id}_{\mathcal{G}}$ as the underlying learning algorithm enables computing of the thresholds and statistics even when the space $\mathcal{G}$ is infinite-dimensional. This last point is the reason why it is essential that Theorem 4.3 applies to infinite-dimensional outputs with non-trace-class kernels.
Together, these explanations show that we can specialize the test of conditional expectations into a test of more complex functionals of Markov kernels — including a conditional two-sample test — while preserving computational feasibility. Indeed, the resulting computational cost is equivalent to that of KRR (with the chosen bootstrapping scheme, if thresholds are boostrapped). We conclude this section with a numerical study highlighting the effectiveness of the test for different functionals.
We now study whether there are many functionals of interest that satisfy Assumption 4.5, and how to pick or find the corresponding map $\Phi_{\mathcal{F}}$. We begin with an example exhibiting a suitable map $\Phi_{\mathcal{F}}$ for the functional $\mathcal{F}$ extracting the conditional expectation and variance, and see how to leverage KMEs of distributions to capture a large class of functionals with explicit maps $\Phi_{\mathcal{F}}$, as it is required in the test procedure. Then, we explain how using KRR with a diagonal kernel together with this idea yields the popular framework of conditional mean embeddings [38], enabling tractable computations even when $\mathcal{G}$ is infinite-dimensional. We conclude this section with a numerical study highlighting how different choices of $\mathcal{F}$ and $\Phi_{\mathcal{F}}$ affect the test outcome.
A.1 Example: conditional expectation and variance
We begin with an example exhibiting a suitable representation map $\Phi_{\mathcal{F}}$ of the functional $\mathcal{F}$ extracting the conditional expectation and variance. Assume that $\mathcal{Z}\subset\mathbb{R}$, and that $p_{1}(\cdot,x)$ and $p_{2}(\cdot,x)$ have well-defined first and second moments for all $x\in\mathcal{X}$. This is for instance the case if $\mathcal{Z}$ is bounded. Introduce for $i\in\{1,2\}$ the first two conditional moments:
$\mu^{(i)}_{1}(x)=\mathbb{E}_{Z\sim p_{i}(\cdot,x)}[Z],\quad\text{and}\quad\mu^{(i)}_{2}(x)=\mathbb{E}_{Z\sim p_{i}(\cdot,x)}[Z^{2}],$
for all $x\in\mathcal{X}$. Now define the functional $\mathcal{F}$ that maps a measure to the vector of its first two moments. For $i\in\{1,2\}$, we have
$\mathcal{F}(p_{i}(\cdot,x))$ $:=\begin{pmatrix}\mu^{(i)}_{1}(x)\\ \mu^{(i)}_{2}(x)\end{pmatrix}$
$=\begin{pmatrix}\mathbb{E}_{Z\sim p_{i}(\cdot,x)}[Z]\\ \mathbb{E}_{Z\sim p_{i}(\cdot,x)}[Z^{2}]\end{pmatrix}$
$=\begin{pmatrix}\int_{\mathcal{Z}}zp_{i}(\mathrm{d}z,x)\\ \int_{\mathcal{Z}}z^{2}p_{i}(\mathrm{d}z,x)\end{pmatrix}$
$=\int_{\mathcal{Z}}\Phi_{\mathcal{F}}(z)p_{i}(\mathrm{d}z,x),$
where we introduced $\Phi_{\mathcal{F}}:z\in\mathcal{Z}\mapsto\begin{pmatrix}z&z^{2}\end{pmatrix}^{\top}$. This shows that the functional $\mathcal{F}$ satisfies Assumption 4.5 with $\mathcal{G}=\mathbb{R}^{2}$.
Let us now spell out the method summarized in Section 4.3 to obtain a test of equality of the first two moments. Assuming that one has access to processes of transition pairs $D^{(i)}=((X_{n}^{(i)},Z_{n}^{(i)}))_{n\in{\mathbb{N}_{>0}}}$, replace them with the processes
$D^{(i)}_{\Phi_{\mathcal{F}}}=((X_{n}^{(i)},Y^{(i)}_{n}))_{n\in{\mathbb{N}_{>0}}}:=((X_{n}^{(i)},\Phi_{\mathcal{F}}(Z_{n}^{(i)})))_{n\in{\mathbb{N}_{>0}}}=\left(\left(X_{n}^{(i)},\begin{pmatrix}Z^{(i)}_{n}\\ {Z^{(i)}_{n}}^{2}\end{pmatrix}\right)\right)_{n\in{\mathbb{N}_{>0}}},$
and use a learning algorithm to learn the map from $(X_{n}^{(i)})_{n}$ to $(Y_{n}^{(i)})_{n}$ for each $i\in\{1,2\}$ (formally, the learning algorithm should seek to approximate the conditional expectation function of $Y_{n}^{(i)}$ conditioned on $X^{(i)}_{n}$, which we assume is independent of $n$). As a consequence of the data augmentation step $y=\Phi_{\mathcal{F}}(z)$, notice that the output of this learning algorithm should be $2$-dimensional. Theorem 4.2 then guarantees that comparing the distance between the predictions of the two models to the sum of confidence bounds on the models’ accuracy provides a test with prescribed level of the hypothesis that the first and second moments of $p_{1}$ and $p_{2}$ coincide.
A.2 Generalization via kernel mean embeddings of distributions
We now generalize the ideas of the above example to find appropriate representation maps $\Phi_{\mathcal{F}}$ of functionals of interest, as the testing procedure requires knowing these maps. The generalization is based on the observation that, for a rich class of functions $\Phi:\mathcal{Z}\to\mathcal{G}$, the integrals involved in Assumption 4.5 are well-studied objects and capture specific properties of the underlying distributions. Furthermore, we argue that any functional $\mathcal{F}$ satisfying Assumption 4.5 has a feature map belonging to the aforementioned class (Proposition A.1), justifying restricting our interest to this class.
Let us begin with this second point and assume that we have a map $\Phi:\mathcal{Z}\to\mathcal{G}$ available, where $\mathcal{G}$ is a Hilbert space, and let $\mathcal{F}$ be a functional of which $\Phi$ is a representation map. Another name for such a $\Phi$ is a feature map on $\mathcal{Z}$, as it can be thought of as extracting information from a point $z\in\mathcal{Z}$ and representing it in $\mathcal{G}$. Readers familiar with RKHS theory may now recognize that the pair $(\Phi,\mathcal{G})$ is a feature map/feature space pair, and it defines a unique RKHS on $\mathcal{Z}$ via the usual canonical isomorphism between a feature space and the associated RKHS [11, Theorem 4.21]. In more practical terms, the function $\Phi$ defines a kernel
$\kappa(z,z^{\prime})=\langle\Phi(z),\Phi(z^{\prime})\rangle_{\mathcal{G}},$
and its RKHS $\mathcal{H}_{\kappa}$ is isometrically isomorphic to the closed span of the family $\{\Phi(z)\mid z\in\mathcal{Z}\}$ in $\mathcal{G}$. It follows immediately that the map
$\Phi_{\kappa}:z\in\mathcal{Z}\mapsto\kappa(\cdot,z)\in\mathcal{H}_{\kappa}$
also is a representation map of the same functional $\mathcal{F}$. Summarizing, it is always possible to pick $\mathcal{G}$ to be \@iaciRKHS RKHS in Assumption 4.5, and $\Phi$ the associated canonical feature map. We formalize this in the following result.
Proposition A.1.
Let $\mathcal{F}$ be a functional satisfying Assumption 4.5, and denote by $\Phi_{0}:\mathcal{Z}\to\mathcal{G}_{0}$ a representation map of $\mathcal{F}$. Let
$\kappa:(z,z^{\prime})\in\mathcal{Z}\mapsto\langle\Phi_{0}(z),\Phi_{0}(z^{\prime})\rangle_{\mathcal{G}_{0}}.$
Then, $\kappa$ is a kernel, and the map
$\Phi_{\kappa}:z\in\mathcal{Z}\mapsto\kappa(\cdot,z)\in\mathcal{H}_{\kappa}$ (13)
also satisfies Assumption 4.5 for $\mathcal{F}$.
Consequences
Given a functional $\mathcal{F}$ satisfying Assumption 4.5, let us take a kernel function $\kappa$ such that $\Phi_{\kappa}$ is a representation map of $\mathcal{F}$. The assumption guarantees that, for any $P,Q\in{\mathcal{M}_{1}^{+}}(\mathcal{Z})$, we have $\mathcal{F}(P)=\mathcal{F}(Q)$ if, and only if,
$\int_{\mathcal{Z}}\kappa(\cdot,z)\mathrm{d}P(z)=\int_{\mathcal{Z}}\kappa(\cdot,z)\mathrm{d}Q(z),$ (14)
where the Bochner integrals are assumed to exist. We can immediately recognize that (14) is an equality between the KMEs of the measures $P$ and $Q$ for the kernel $\kappa$. Introducing the KME map
$\Pi_{\kappa}:M\in{\mathcal{M}_{1}^{+}}(\mathcal{Z})\mapsto\int_{\mathcal{Z}}\kappa(\cdot,z)\mathrm{d}M(z),$
the condition (14) simply rewrites as $\Pi_{\kappa}(P)=\Pi_{\kappa}(Q)$. This observation is essential in practice, as the question of what information about a distribution is captured by its KME is a well-studied area of research [38]. For instance, the inhomogeneous polynomial kernel of order $m\in{\mathbb{N}_{>0}}$ captures the first $m$ moments of a distribution; the exponential kernel captures the moment-generating function; and the Gaussian kernel captures the full distribution as it is characteristic [38, Sections 3.1.1 and 3.3.1]. In other words — and this is the concluding observation of this section:
Observation A.2 (Picking $\Phi$ in Assumption 4.5).
When the map $\Phi$ is chosen as the partial evaluation of a kernel function $\kappa$ as in (13), the corresponding hypothesis (12) asserts the equality of the corresponding KMEs of the conditional distributions. Therefore, the properties being tested are directly the ones these KMEs capture, and they depend on the choice of the kernel function $\kappa$.
A.3 The special case of KRR with a diagonal kernel: CMEs and their implementation
The challenge when following Observation A.2 and picking the representation map to be a kernel canonical feature map is that the data is now composed of functions; specifically, the data sets have the form $((X_{n}^{(i)},\kappa(\cdot,Z_{n}^{(i)})))_{n}$, and the model output should be an element of $\mathcal{H}_{\kappa}$. This may be impossible to handle for some learning methods; e.g., feedforward neural networks, which need an output of fixed, finite dimension. Furthermore, even when it is possible to train and evaluate the algorithm with such data, computing the test statistic or threshold in (7) may still reveal intractable.
Fortunately, both of these concerns are void for KRR with a diagonal kernel $K=k\cdot\mathrm{id}_{\mathcal{H}_{\kappa}}$ (with $k$ a scalar kernel): it accepts functional data seamlessly, and the computations are tractable. As a matter of fact, KRR with a diagonal kernel and such data recovers exactly the popular framework of CMEs [39].
Before proceeding, we emphasize that we now have two distinct kernel functions: the kernel $K$, which is used to perform KRR, and the kernel $\kappa$, which is chosen depending on the functional we wish to test for in (12) accordingly to the developments of the previous section. The two are connected by the relation $K=k\cdot\mathrm{id}_{\mathcal{H}_{\kappa}}$, where $k$ is a freely-chosen scalar kernel.
Computation of statistic and thresholds
We are interested in evaluating the statistic and threshold involved in (7). The computations to achieve this are commonly known [38, Section 4.1.2], and we report them here for completeness.
A first observation is that the quantities involved in the threshold — specifically, those defined in ˜8, 9, 11 and 10 — only involve the kernel $K$. This kernel indirectly depends on $\kappa$ since $K$ takes values in $\mathcal{L}_{\mathrm{b}}(\mathcal{H}_{\kappa})$, but we assume that we are able to compute the quantities involved for $K$. In particular, under the UBD assumption, this assumption directly translates to the feasibility of these computations for $K_{0}$}, which is trivially satisfied if $\tilde{\mathcal{G}}$ is finite-dimensional. This is the case in the standard case where $K=k\cdot\mathrm{id}_{\mathcal{H}_{\kappa}}$, since $\tilde{\mathcal{G}}$ is simply $\mathbb{R}$.
Next is the computation of the statistic, $\lVert f_{D_{1},\lambda_{1}}(x)-f_{D_{2},\lambda_{2}}(x)\rVert_{\mathcal{H}_{\kappa}}$, $x\in\mathcal{X}$. Importantly, for every $x\in\mathcal{X}$, ignoring the subscripts 1 and 2 indexing over the two processes at hand for readability,
@@ A.3 The special case of KRR with a diagonal kernel: CMEs and their implementation
$f_{D,\lambda}(x)=K(x,X_{:N})(K(X_{:N},X_{:N})+\lambda\mathrm{id}_{\mathcal{H}_{\kappa}^{N}})^{-1}\begin{pmatrix}\kappa(\cdot,Z_{1})\\ \vdots\\ \kappa(\cdot,Z_{N})\end{pmatrix}$.
We now leverage the fact that $K=k\cdot\mathrm{id}_{\mathcal{H}_{\kappa}}$. The identity operator on $\mathcal{H}_{\kappa}$ factors out and we obtain
$f_{D,\lambda}(x)=\left(z\in\mathcal{Z}\mapsto k(x,X_{:N})(k(X_{:N},X_{:N})+\lambda\mathrm{id}_{\mathbb{R}^{N}})^{-1}\begin{pmatrix}\kappa(z,Z_{1})\\ \vdots\\ \kappa(z,Z_{N})\end{pmatrix}\right)$.
Introducing now the coefficients
$\alpha^{(i)}_{n}(x)=(k(X_{:N}^{(i)},X_{:N}^{(i)})+\lambda_{i}\mathrm{id}_{\mathbb{R}^{N}})^{-1}k(X_{:N}^{(i)},x),$
for all $i\in\{1,2\}$ and with $N=\mathrm{len}(D_{i})$, we obtain
$\lVert f_{D_{1},\lambda_{1}}(x)-f_{D_{2},\lambda_{2}}(x)\rVert_{\mathcal{H}_{\kappa}}^{2}$
$\quad=\sum_{n,m=1}^{\mathrm{len}(D_{1})}\alpha_{n}^{(1)}(x)\alpha_{m}^{(1)}(x)\kappa(Z_{n}^{(1)},Z_{m}^{(1)})+\sum_{n,m=1}^{\mathrm{len}(D_{2})}\alpha_{n}^{(2)}(x)\alpha_{m}^{(2)}(x)\kappa(Z_{n}^{(2)},Z_{m}^{(2)})$
$\qquad\qquad-2\sum_{n=1}^{\mathrm{len}(D_{1})}\sum_{m=1}^{\mathrm{len}(D_{2})}\alpha_{n}^{(1)}(x)\alpha_{m}^{(2)}(x)\kappa(Z_{n}^{(1)},Z_{m}^{(2)}).$
Here, we used the identity $\kappa(z,z^{\prime})=\langle\kappa(\cdot,z),\kappa(\cdot,z^{\prime})\rangle_{\mathcal{H}_{\kappa}}$, for all $z,z^{\prime}\in\mathcal{Z}$. Concluding, if the underlying learning algorithm is KRR with kernel $K=k\cdot\mathrm{id}_{\mathcal{H}_{\kappa}}$, it is possible to evaluate the statistic and threshold of the test of Theorem 4.4. We provide in Appendix C algorithms that summarize these computations for implementation purposes.
A.3.1 Numerical study